OpenStudy (danysanchy97):

Juan Ramirez sells suits in a major department store on weekends. He earns a commission of 5 percent on the first ten suits, and if he sells more than ten, he earns an additional 3 percent on the additional suits. Last weekend Juan sold thirteen $250 suits. What was his commission? A. $210.40 B. $260.00 C. $147.50 D. $185.00

OpenStudy (phi):

He earns a commission of 5 percent on the first ten suits, First, find how much does it cost to buy 10 suits. Any idea ?

OpenStudy (danysanchy97):

$2,500??

OpenStudy (phi):

yes, 10*250 = 2500 next, find 5% of 2500 that means change 5% to 5/100 and figure out \[ \frac{5}{100} \cdot 2500 \]

OpenStudy (danysanchy97):

125?

OpenStudy (phi):

yes. now we need to find the commission on the rest of the suits. (3 more, right?)

OpenStudy (phi):

it says he gets 5 percent plus an additional 3 percent on the next 3 suits what percent is that ?

OpenStudy (phi):

5% + 3% is 8% he gets 8% of 3 suits in other words, he gets 8% of 3*250 can you figure that out ?

OpenStudy (phi):

8% of 3*250 first, change 8% to 8/100 and figure out \[ \frac{8}{100} \cdot 3 \cdot 250 \]